SERIAL NUMBERS FROM A PRODUCT ARE TO BE MADE USING 2 LETTERS FOLLOWED BY 3 NUMBERS IF THE LETTERS ARE TAKEN FROM THE FIRST 6 LETTERS OF THE ALPHABET WITH NO REPEATS AND THE NUMBERS ARE TAKEN FROM 0 TO 9 WITH NO REPEATS HOW MANY SERIAL NUMBERS ARE POSSIBLE
Answer by math_helper(2461)


The first character can be made in 6 ways
The 2nd in 5 ways (because no repeats)
For the 3rd character there are 10 choices
4th: 9 (because no repeats, there are only 9 choices)
5th: 8 (because no repeats, there are only 8 choices)
Total number of serial numbers = (6*5)*(10*9*8) = (30)*(720) = 21600